# Random Transverse Field Effects on Magnetic Noise in Spin Systems

### v1

@@ F.2 Random-walk simulations
To obtain the results presented in Fig. 2, we simulate the random walk developed in Sec. III.2 and App. C. We initialize a system with $6000$ spins and periodic boundary condition and construct the corresponding RW-lattice on which our simulation run; a cutoff cluster size $n=10$ is used. A random walker is initialized randomly across the RW-lattice sites; in each step, the random walker moves with probability $w\delta t$ to one of its adjacent sites, where $w$ is the corresponding transition rate. We select $\delta t=1/50$. The data presented in Fig. 2b are averaged over $200$ realizations of disorders.

### v2

@@ F.2 Random-walk simulations
To obtain the results presented in Fig. 2, we simulate the random walk developed in Sec. III.2 and App. C. We initialize a system with $6000$ spins and periodic boundary condition and construct the corresponding random walk lattice on which our simulation run; a cutoff cluster size $n=10$ is used. A random walker is initialized randomly across the random walk lattice sites; in each step, the random walker moves with probability $w\delta t$ to one of its adjacent sites, where $w$ is the corresponding transition rate. We select $\delta t=1/50$. The data presented in Fig. 2b are averaged over $200$ realizations of disorders.